# PartComposer: Learning and Composing Part-Level Concepts from Single-Image Examples

### v1

@@ Abstract.
We present PartComposer: a framework for part-level concept learning from single-image examples that enables text-to-image diffusion models to compose novel objects from meaningful components. Existing methods either struggle with effectively learning fine-grained concepts or require a large dataset as input. We propose a dynamic data synthesis pipeline generating diverse part compositions to address one-shot data scarcity. Most importantly, we propose to maximize the mutual information between denoised latents and structured concept codes via a concept predictor, enabling direct regulation on concept disentanglement and re-composition supervision. Our method achieves strong disentanglement and controllable composition, outperforming subject and part-level baselines when mixing concepts from the same, or different, object categories.
personalization, multiple visual concept extraction, information theory, example-based modeling
ccs: Computing methodologies Machine learning; Computer graphics
Figure 1. Given one-shot images that contains multiple fine-grained parts, our pipeline can learn descriptive concepts for these parts with disentanglement and flexibly re-compose them to generate new objects, for both intra-category and cross-category objects.
@@ 1. Introduction
Visually inspired and creative generation emerges from the ability to compose new objects from familiar parts (Richardson et al., 2024, 2025). From virtual creatures to fantastical designs, part-level concept composition is a powerful paradigm for visual imagination. For instance, given images of several example chairs (Figure 1), designers could imagine composing a new chair that contains parts borrowed across the example chairs and is structurally realistic. However, there are two major challenges for this task. Firstly, the part-level concepts are more fine-grained than subject-level concepts, and require structural information to be composed into reasonable objects. This makes it more challenging to properly learn the identity of concepts and achieve clear disentanglement between different concepts. Secondly, without dedicated 3D models or multi-view images of objects, most real-world examples in concept learning and composition are single-image examples (i.e., there is just one image per object). Such scarce data results in very poor data variety (i.e., very limited ground truth part composition), challenging the capability of generative models to properly compose unseen combinations.
@@
A commonly used approach in concept learning is to extract visual concepts from generative models in the form of latent codes (Gal et al., 2022; Ruiz et al., 2023; Kumari et al., 2023; Han et al., 2023; Wei et al., 2023; Jia et al., 2023; Shi et al., 2024). Each of the extracted concepts encodes the identity of an image object (e.g., a red seat cushion) and can be used with other concepts in creative image generation (Avrahami et al., 2023; Ng et al., 2024; Richardson et al., 2024; Jang et al., 2024). In this context, text-to-image diffusion models serve as a versatile tool for learning compositional concepts through personalization. A common paradigm is to learn new or specialized token embeddings for concepts, through finetuning the diffusion model (updating the embeddings and/or the model weights) via latent diffusion loss. The learned concepts token embeddings can be directly used in prompts at inference time to generate images containing the target concepts. Several recent methods have explored dealing with single-image inputs (Wei et al., 2023; Jia et al., 2023; Shi et al., 2024; Avrahami et al., 2023; Garibi et al., 2025), but they operate at the subject level. Other methods focus on part-level concept learning (Ng et al., 2024; Richardson et al., 2025), but they require a dataset of images as input. A common problem is that these methods tend to fail to disentangle and retain the identity of fine-grained parts. Although several works (Avrahami et al., 2023; Ng et al., 2024) propose to use cross-attention loss between concept tokens and the visual contents to disentangle concepts, our experiments show that this method alone cannot effectively deal with part-level concepts from one-shot inputs. Figure 2 illustrates this problem when composing 4 parts across 2 chairs, where some part-levels concepts are ambiguous or even missing in composing new part combinations.
@@
Figure 2. Illustration of the challenge in learning fine-grained concepts from one-shot inputs. Break-a-Scene (Avrahami et al., 2023) results in multiple entanglement while out method cleanly reflect target concepts.
@@
To overcome these challenges, we make two critical observations and propose corresponding modifications. Firstly, proper augmentation on the possible part combinations should be explored and the structural integrity of objects should be preserved in the generative model prior. We propose a dynamic data synthesis pipeline to generate rich part-level supervision from single-image examples, by grouping union-sampled input images with randomly synthesized part combination images. Secondly, we observe that there is lack of regulation of the information encoded in different concepts, resulting in entangled and ambiguous concept learning and composition. We propose a mutual information maximization framework that explicitly aligns denoised latents with structured part-level concept codes, by inferring which concepts are preserved given denoised latents. To effectively formalize such regulation, we argue that penalizing both wrong compositions and wrong localization of concepts are essential to achieve good concept disentanglement. Thus, we introduce a concept predictor that performs both classification and segmentation on latent features, regularizing the embedding space to reflect concept presence and spatial structure.
@@
We evaluate our methods on single-image part composition tasks from same object categories and across different categories. We compare the part composition results with several baselines (Avrahami et al., 2023; Ng et al., 2024; Jang et al., 2024; Richardson et al., 2025) in terms of concept preserving capability and visual quality. Our method significantly out-performs baselines in preserving correct identity of intended part-level concepts, while maintaining image quality. When given cross-category examples, our methods can produce interesting part composition of virtual objects with reasonable structure (Figure 1), enabling great artistic imagination. We also provide an ablation study for our key designs and demonstrate the flexibility of our methods on different input forms.
@@
In summary, our contributions are: (1) We propose to focus on a challenging task, learning part-level concepts from single-image inputs, that relaxes input requirement and can be directly applied on most real-world images. (2) We propose a dynamic data synthesis pipeline to automatically augment the scarce one-shot inputs and balance between structural integrity and part combination variety. (3) We propose a maximizing information scheme to explicitly disentangle the information encoded in concepts and supervise composition through a concept predictor with classification and segmentation.
@@ 2. Background in Visual Concept Learning
Pioneering works like Textural Inversion (Gal et al., 2022) and DreamBooth (Ruiz et al., 2023) in visual concept learning require multiple images as input to encode a single a concept. Several follow-up works have loosen the input constraint as learning single concept from single-image examples (Wei et al., 2023; Jia et al., 2023; Shi et al., 2024). Multi-concept learning from multiple inputs are also being discovered (Kumari et al., 2023; Han et al., 2023). Break-a-Scene (Avrahami et al., 2023) first propose a pipeline to learn multiple subject-level concepts from single images, enabling great flexibility in learning and remixing concepts from general use cases.
@@
Most concept learning methods focus on learning concept for the entire image or on the subject level (Wei et al., 2023; Jia et al., 2023; Shi et al., 2024; Kumari et al., 2023; Han et al., 2023; Avrahami et al., 2023; Jin et al., 2024; Hao et al., 2024; Jang et al., 2024; Garibi et al., 2025). The target use case for these methods is to generate images that are organically composed by required subject(s) according to user-specified prompts. However, to enable more creativity in generating new subjects, part-level concept learning is required. This more fined-grained concept learning task has been explored by PartCraft (Ng et al., 2024). They learn a large dictionary of concepts for different parts of creatures and generate new virtual creatures that have not been seen in the real world. However, their method rely on training on a large dataset and struggle on single-image examples. Piece-it-Together (PiT) (Richardson et al., 2025) also targets part-level concept learning, by directly operating in a carefully chosen IP-Adapter+ (Ye et al., 2023) representation space and synthesizes a complete and coherent concept by training a generative model to fill in missing information conditioned on a strong domain-specific prior. Unlike optimization-heavy approaches, PiT enables efficient inference, supports diverse sampling from sparse inputs, and allows flexible semantic manipulation. However, PiT requires training on class-specific datasets, where single-image examples are not supported and generalization to out-of-distribution data is limited.
@@
A common challenge in multi-concept learning is to disentangle the identity of concepts (Avrahami et al., 2023; Ng et al., 2024; Jang et al., 2024). To achieve good disentanglement, dynamic masking different compositions of concepts and using cross-attention mechanisms to regulate the diffusion model finetuning are commonly used in methods like Break-a-Scene (Avrahami et al., 2023) and PartCraft (Ng et al., 2024). MuDI (Jang et al., 2024) proposes a dynamic concept composition method and mean-shifted inference technique to further improve the decoupling of different concepts. However, concepts missing or inaccurate identity still occur in all of these methods when re-composing concepts in generative process (Avrahami et al., 2023; Ng et al., 2024; Jang et al., 2024), especially when dealing with 4 or more concepts. Our observation is that all existing concept learning works do not provide any explicit regulation on the information encoded in different concepts and the composition process, causing inaccurate multi-concept remixing. Other common challenges include learning large number of concepts (e.g., more than 4) from single inputs (Avrahami et al., 2023).
@@ 3. Method
Given single-image examples, we want to be able to learn part concepts and re-compose them from each input together arbitrarily. Our pipeline build upon the commonly used diffusion model customization approaches (introduced in Section 1) like Break-a-Scene (Avrahami et al., 2023) and PartCraft (Ng et al., 2024), where standard diffusion loss $\mathcal{L_{\text{ldm}}}$ and cross-attention losses $\mathcal{L_{\text{attn}}}$ are used to learn and disentangle concept tokens. However, to enable learning concepts at part level with single-image examples, we propose a dynamic data synthesis method to augment the limited data and a maximizing mutual information scheme to enable clear disentanglement of parts’ concepts and good composition capabilities for the generative model. Figure 4 gives an overview of our approach.
We want to highlight that although our pipeline is built on text-to-image diffusion models, we treat part-level concepts as purely visual information and do not require semantic or annotations for the part concepts. The tokens and text prompts only serve as media of templates to encode and convey such information.
@@
Figure 4. Overview of our method. Given a concept-compositional prompt and noise input, the denoising U-Net produces a latent $\tilde{z}$ supervised by reconstruction loss $\mathcal{L_{\text{ldm}}}$, cross-attention losses $\mathcal{L_{\text{attn}}}$, and information loss $\mathcal{L_{\text{info}}}$. $\mathcal{L_{\text{info}}}$ is computed by a concept predictor which receives $\tilde{z}$ and outputs concept classification and segmentation logits. The goal is to maximize mutual information between latent features and concept codes. All modules are jointly optimized to enable part-level concept disentanglement and controllable composition.
@@
We propose a dynamic data synthesis approach to augment the limited single-image example input. Our goal is to explore and augment part combination variety while preserving structural integrity information in the diffusion prior. We formalize the training batch as two images, an instance image that is directly sampled from the given examples and a synthetic image that is generated on-the-fly. Figure 3 demonstrates our dynamic data synthesis approach. To preserve the structural information of the diffusion prior, the instance image is randomly chosen from the input examples, since they contain reasonable object structures. We randomly select a subset of the parts contained in the chosen instance image, inspired by the union sampling method in Break-a-Scene (Avrahami et al., 2023). We mask out the unused parts areas and change the background to white to focus only on the part concepts, and pair it with a descriptive prompt. However, this data alone cannot provide enough training data to finetune diffusion models to compose different compositions of parts across different images. Thus, to explore and augment part combination variety, we propose to synthesize an image that contains parts that are randomly sampled across input examples. For each part category (e.g., the armrest of the chair), we randomly select a part from the input instances. Inspired by MuDI (Jang et al., 2024), we randomly scale and place the sampled parts on an white image. The parts can overlap each other to encourage learning concepts from multiple possible compositions. We modify the original masks according to the overlapping occlusions. Due to the high variability in the synthetic image, we use a different prompt to treat it as a collection of parts. Additional dynamic data synthesis explanation and examples can be found in Appendix A.
@@
Most existing concept learning methods either rely on multiple inputs (Kumari et al., 2023; Han et al., 2023; Ng et al., 2024) or are restricted by the number of concepts they could learn at the same time (Wei et al., 2023; Jia et al., 2023; Shi et al., 2024). These restrictions results in poor performance in applying them on learning part-level concepts. We argue that a key missing point is that there are no explicit regulation on the information encoded in the concept embeddings and the composition process of the generative model. We thus propose a scheme to maximize the mutual information between the denoised latents and the concepts contained in the input image, and thus disentangle the information encoded in each part-level concepts and supervise the diffusion generative process.
To this end, we adopt a mutual information maximization objective inspired by InfoGAN (Chen et al., 2016), which encourages the learned concept embeddings to retain interpretable and disentangled semantic structure. Let $\tilde{z}$ denote the denoised latent representation, and let $\mathbf{c}$ denote the set of concept codes associated with the input image. We seek to maximize the mutual information $I(\mathbf{c};\tilde{z})$, which quantifies how much information about the concepts is preserved in the latent. This can be expressed as:
$I(\mathbf{c};\tilde{z})=H(\mathbf{c})-H(\mathbf{c}\mid\tilde{z}),$
where $H(\mathbf{c})$ is the entropy of the concept distribution and $H(\mathbf{c}\mid\tilde{z})$ is the conditional entropy of the concepts given the latent. Since directly computing this term is intractable, we introduce a variational distribution $Q(\mathbf{c}\mid\tilde{z})$, implemented via a neural network - concept predictor, to approximate the true posterior. This leads to a variational lower bound:
@@
which satisfies $\mathcal{I}_{\text{lower}}\leq I(\mathbf{c};\tilde{z})$. For a fixed prior over concepts (i.e., the ground-truth labels that supervise the concept predictor), $H(\mathbf{c})$ is constant and can be omitted during optimization. The corresponding training loss $\mathcal{L}_{\text{Info}}$ is defined as the negative of the lower bound :
$\mathcal{L}_{\text{Info}}=-\mathbb{E}_{\mathbf{c}\sim P(\mathbf{c}),\tilde{z}\sim P(\tilde{z}\mid\mathbf{c})}[\log Q(\mathbf{c}\mid\tilde{z})],$
which is minimized alongside the task loss. This encourages the model to produce latents $\tilde{z}$ from which the concept codes $\mathbf{c}$ can be accurately inferred, effectively regularizing the concept embedding space to reflect correct part-level visual information. This framework provides direct supervision to disentangle part-level concepts and penalizes the generative models from generating ambiguous or missing part combinations.
To effectively implement a $Q(\mathbf{c}\mid\tilde{z})$, we propose the concept predictor design as shown in Figure 4, which includes two output heads to provide both classification and segmentation predictions of concepts given an denoised latent. The classification loss $\mathcal{L_{\text{CLS}}}$ penalizes the wrong compositions of concepts (i.e., missing concepts or containing more concepts). The segmentation loss $\mathcal{L_{\text{SEG}}}$ further penalizes wrong localization of concepts (i.e., wrong or entangled location of concepts). An visualized example of the ground truth localization masks and predicted localization segmentation are shown in Figure 4. We also provide visualization for convergence of the segmentation head throughout the training steps in Appendix D. These two losses are weighted to have the same scale and combined to get the mutual information loss $\mathcal{L}_{\text{Info}}$. The concept predictor is jointly optimized with the concept learning process.
@@
Our overall training loss consists the latent diffusion loss $\mathcal{L}_{\text{ldm}}$ and cross-attention loss $\mathcal{L}_{\text{attn}}$ from the original Break-a-Scene pipeline, with our concept prediction loss composed by the classification loss $\mathcal{L}_{\text{CLS}}$ and segmentation loss $\mathcal{L}_{\text{SEG}}$. We also use an auxiliary background loss $\mathcal{L}_{\text{BG}}$ to further improve the generated image quality, which penalizes generation of content outside the union of selected part masks. This encourages the model to focus on concept-relevant regions and avoid unrelated artifacts in unmasked areas. Thus, the total loss is defined as: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{ldm}}+\mathcal{L}_{\text{attn}}+\lambda_{\text{info}}(\mathcal{L}_{\text{CLS}}+\lambda_{\text{seg}}\mathcal{L}_{\text{SEG}})+\lambda_{\text{bg}}\mathcal{L}_{\text{BG}},$ with the following loss weights: $\lambda_{\text{info}}=0.05$, $\lambda_{\text{seg}}=10.0$, $\lambda_{\text{bg}}=0.01$.

### v2

@@ Abstract.
We present PartComposer: a framework for part-level concept learning from single-image examples that enables text-to-image diffusion models to compose novel objects from meaningful components. Existing methods either struggle with effectively learning fine-grained concepts or require a large dataset as input. We propose a dynamic data synthesis pipeline generating diverse part compositions to address one-shot data scarcity. Most importantly, we propose to maximize the mutual information between denoised latents and structured concept codes via a concept predictor, enabling direct regulation on concept disentanglement and re-composition supervision. Our method achieves strong disentanglement and controllable composition, outperforming subject and part-level baselines when mixing concepts from the same, or different, object categories. Our code is released in https://github.com/Junyu-Liu-Nate/partcomposer.
personalization, multiple concept extraction, information theory
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3757377.3763947conference: SIGGRAPH Asia’ 25 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongisbn: 979-8-4007-2137-3/2025/12submissionid: 139ccs: Computing methodologies Machine learning; Computer graphics
Figure 1. Given single-image examples where each image contains multiple fine-grained parts, our pipeline can learn descriptive concepts for these parts with clear disentanglement and flexibly re-compose them to generate new objects, for both intra-category and cross-category objects.
@@ 1. Introduction
Visually inspired and creative generation emerges from the ability to compose new objects from familiar parts (Richardson et al., 2024, 2025). From virtual creatures to fantastical designs, part-level concept composition is a powerful paradigm for visual imagination. For instance, given images of several example chairs (Fig. 1), designers could imagine composing a new chair that contains parts borrowed across the example chairs and is structurally realistic. However, there are two major challenges for this task. Firstly, the part-level concepts are more fine-grained than subject-level concepts, and require structural information to be composed into reasonable objects. This makes it more challenging to properly learn the identity of concepts and achieve clear disentanglement between different concepts. Secondly, without dedicated 3D models or multi-view images of objects, most real-world examples in concept learning and composition are single-image examples (i.e., there is just one image per object). Such scarce data results in very poor data variety (i.e., very limited ground truth part composition), challenging the capability of generative models to properly compose unseen combinations.
@@
A commonly used approach in concept learning is to extract visual concepts from generative models in the form of latent codes (Gal et al., 2022; Ruiz et al., 2023; Kumari et al., 2023; Han et al., 2023; Wei et al., 2023; Jia et al., 2023; Shi et al., 2024). Each of the extracted concepts encodes the identity of an image object (e.g., a red seat cushion) and can be used with other concepts in creative image generation (Avrahami et al., 2023; Ng et al., 2024; Richardson et al., 2024; Jang et al., 2024). In this context, text-to-image diffusion models serve as a versatile tool for learning compositional concepts through personalization. A common paradigm is to learn new or specialized token embeddings for concepts, through fine-tuning the diffusion model (updating the embeddings and/or the model weights) via latent diffusion loss. The learned concept token embeddings can be directly used in prompts at inference time to generate images containing the target concepts. Several recent methods have explored dealing with single-image inputs (Wei et al., 2023; Jia et al., 2023; Shi et al., 2024; Avrahami et al., 2023; Garibi et al., 2025), but they operate at the subject level. Other methods focus on part-level concept learning (Ng et al., 2024; Richardson et al., 2025), but they require a dataset of images as input. A common problem is that these methods tend to fail to disentangle and retain the identity of fine-grained parts. Although several works (Avrahami et al., 2023; Ng et al., 2024) propose to use cross-attention loss between concept tokens and the visual contents to disentangle concepts, our experiments show that this method alone cannot effectively deal with part-level concepts from one-shot inputs. Fig. 2 illustrates this problem when composing 4 parts across 2 chairs, where some part-level concepts are ambiguous or even missing in composing new part combinations.
@@
Figure 2. Illustration of the challenge in learning fine-grained concepts from single-image examples. Break-a-Scene (Avrahami et al., 2023) results in multiple entanglements while our method cleanly reflects target concepts.
@@
To overcome these challenges, we make two critical observations and propose corresponding modifications. Firstly, proper augmentation on the possible part combinations should be explored, and the structural integrity of objects should be preserved in the generative model prior. We propose a dynamic data synthesis pipeline to generate rich part-level supervision from single-image examples, by grouping union-sampled input images with randomly synthesized part combination images. Secondly, we observe that there is a lack of regulation of the information encoded in different concepts, resulting in entangled and ambiguous concept learning and composition. We propose a mutual information maximization framework that explicitly aligns denoised latents with structured part-level concept codes, by inferring which concepts are preserved given denoised latents. To effectively formalize such regulation, we argue that penalizing both wrong compositions and wrong localization of concepts are essential to achieve good concept disentanglement. Thus, we introduce a concept predictor that performs both classification and segmentation on latent features, regularizing the embedding space to reflect concept presence and spatial structure.
@@
We evaluate our methods on single-image part composition tasks from the same object categories and across different categories. We compare the part composition results with several baselines (Avrahami et al., 2023; Ng et al., 2024; Jang et al., 2024; Richardson et al., 2025) in terms of concept preserving capability and visual quality. Our method significantly out-performs baselines in preserving the correct identity of intended part-level concepts, while maintaining image quality. When given cross-category examples, our methods can produce interesting part composition of virtual objects with reasonable structure (Fig. 1), enabling great artistic imagination. We also provide an ablation study for our key designs and demonstrate the flexibility of our methods on different input forms.
@@
In summary, our contributions are: (1) We propose to focus on a challenging task, learning part-level concepts from single-image inputs, which relaxes the input requirement for objects’ images. (2) We propose a dynamic data synthesis pipeline to automatically augment the scarce one-shot inputs and balance between structural integrity and part combination variety. (3) We propose a maximizing information scheme to explicitly disentangle the information encoded in concepts and supervise composition through a concept predictor with classification and segmentation.
@@ 2. Background in Visual Concept Learning
Pioneering works like Textural Inversion (Gal et al., 2022) and DreamBooth (Ruiz et al., 2023) in visual concept learning require multiple images as input to encode a single concept. Several follow-up works have loosened the input constraint as learning a single concept from single-image examples (Wei et al., 2023; Jia et al., 2023; Shi et al., 2024). Multi-concept learning methods from multiple inputs are also being discovered (Kumari et al., 2023; Han et al., 2023). Break-a-Scene (Avrahami et al., 2023) first proposes a pipeline to learn multiple subject-level concepts from single images, enabling great flexibility in re-compose the learned concepts.
@@
Most concept learning methods focus on learning concepts for the entire image or on the subject level (Wei et al., 2023; Jia et al., 2023; Shi et al., 2024; Kumari et al., 2023; Han et al., 2023; Avrahami et al., 2023; Jin et al., 2024; Hao et al., 2024; Jang et al., 2024; Garibi et al., 2025). The target use case for these methods is to generate images that are organically composed by the required subject(s) according to user-specified prompts. However, to enable more creativity in generating new subjects, part-level concept learning is required. This more fine-grained concept learning task has been explored by PartCraft (Ng et al., 2024). They learn a large dictionary of concepts for different parts of creatures and generate new virtual creatures that have not been seen in the real world. However, their method relies on training on a large dataset and struggles with single-image examples. Piece-it-Together (PiT) (Richardson et al., 2025) also targets part-level concept learning, by directly operating in a carefully chosen IP-Adapter+ (Ye et al., 2023) representation space and synthesizes a complete and coherent concept by training a generative model to fill in missing information conditioned on a strong domain-specific prior. Unlike optimization-heavy approaches, PiT enables efficient inference, supports diverse sampling from sparse inputs, and allows flexible semantic manipulation. However, PiT requires training on class-specific datasets, where single-image examples are not supported and generalization to out-of-distribution data is limited.
@@
A common challenge in multi-concept learning is to disentangle the identity of concepts (Avrahami et al., 2023; Ng et al., 2024; Jang et al., 2024). To achieve good disentanglement, dynamic masking of different compositions of concepts and using cross-attention mechanisms to regulate the diffusion model fine-tuning are commonly used in methods like Break-a-Scene (Avrahami et al., 2023) and PartCraft (Ng et al., 2024). MuDI (Jang et al., 2024) proposes a dynamic concept composition method and mean-shifted inference technique to improve the decoupling of different concepts. However, concepts missing or inaccurate identity still occur in all of these methods when re-composing concepts in the generative process (Avrahami et al., 2023; Ng et al., 2024; Jang et al., 2024), especially when dealing with 4 or more concepts. Our observation is that all existing works do not provide any explicit regulation on the information encoded in different concepts and the composition process, causing inaccurate multi-concept remixing. Other common challenges include learning a large number of concepts (e.g., more than 4) from single inputs (Avrahami et al., 2023).
@@ 3. Method
Given single-image examples, we aim to learn part concepts and arbitrarily re-compose them from each input to generate new objects. Our pipeline build upon the commonly used diffusion model customization approaches (introduced in Section 1) like Break-a-Scene (Avrahami et al., 2023) and PartCraft (Ng et al., 2024), where standard diffusion loss $\mathcal{L_{\text{ldm}}}$ and cross-attention losses $\mathcal{L_{\text{attn}}}$ are used to learn and disentangle concept tokens. However, to tackle the challenges of learning part-level concepts from single-image examples, we need to deal with the extreme data scarcity and provide proper supervision on part-level information. To this end, we propose a dynamic data synthesis method to augment the limited data, and a maximizing mutual information scheme to enable clear disentanglement of parts’ concepts and good composition capabilities for the generative model. Fig. 4 gives an overview of our approach.
We would like to emphasize that, although our pipeline is based on text-to-image diffusion models, we treat part-level concepts as purely visual information and do not require semantics or annotations for these part concepts. The tokens and text prompts only serve as media of templates to encode and convey such information.
@@
Figure 4. Overview of our method. Given a concept-compositional prompt and noise input, the denoising U-Net produces a latent $\tilde{z}$ supervised by reconstruction loss $\mathcal{L_{\text{ldm}}}$, cross-attention loss $\mathcal{L_{\text{attn}}}$, and information loss $\mathcal{L_{\text{info}}}$. $\mathcal{L_{\text{info}}}$ is computed by a concept predictor which receives $\tilde{z}$ and outputs concept classification and segmentation predictions. The goal is to maximize mutual information between the denoised latent features and the intended concept codes. All modules are jointly optimized to enable part-level concept disentanglement and controllable composition without concept missing or ambiguity.
@@
We propose a dynamic data synthesis approach to augment the limited single-image example input. Our goal is to explore and augment the part combination variety while preserving structural integrity information in the diffusion prior. We formalize the training batch as two images, an instance image that is directly sampled from the given examples, and a synthetic image that is generated on-the-fly. Fig. 3 demonstrates our dynamic data synthesis approach. To preserve the structural information of the diffusion prior, we randomly choose an image from the input examples as the instance image, since the input examples already contain reasonable object structures. We randomly select a subset of the parts contained in the chosen instance image, inspired by the union sampling method in Break-a-Scene (Avrahami et al., 2023). We mask out the unused parts areas and change the background to white to focus only on the part concepts, and pair it with a descriptive prompt. However, this data alone cannot provide enough training data to fine-tune diffusion models to compose different compositions of parts across different images. Thus, to explore and augment part combination variety, we propose to synthesize an image that contains parts that are randomly sampled across input examples. For each part category (e.g., the armrest of the chair), we randomly select a part from the input instances. Inspired by MuDI (Jang et al., 2024), we randomly scale and place the sampled parts on a white image. The parts can overlap with each other to encourage learning concepts from multiple possible compositions. We modify the original masks according to the overlapping occlusions. Due to the high variability in the synthetic image, we use a different prompt to treat it as a collection of parts.
@@
When observing the poor performance of existing concept learning methods on learning part-level concepts, we argue that a key missing point is that there are no explicit regulations on the information encoded in the concept embeddings and the composition process of the generative model. We thus propose a scheme to maximize the mutual information between the denoised latents and the concepts contained in the input image, and thus disentangle the information encoded in each part-level concept and supervise the diffusion generative process.
[…]
@@
[…]
which is minimized alongside the overall training loss. This encourages the model to produce latents $\tilde{z}$ from which the concept codes $\mathbf{c}$ can be accurately inferred, effectively regularizing the concept embedding space to reflect correct part-level visual information. This framework provides direct supervision to disentangle part-level concepts and penalizes the generative models from generating ambiguous or missing part combinations.
[…]
@@
Our overall training loss consists of the latent diffusion loss $\mathcal{L}_{\text{ldm}}$ and cross-attention loss $\mathcal{L}_{\text{attn}}$ from the original Break-a-Scene pipeline, with our concept prediction loss composed of the classification loss $\mathcal{L}_{\text{CLS}}$ and segmentation loss $\mathcal{L}_{\text{SEG}}$. We also use an auxiliary background loss $\mathcal{L}_{\text{BG}}$ to further improve the generated image quality, which penalizes the generation of content outside the union of selected part masks. This encourages the model to focus on concept-relevant regions and avoid unrelated artifacts in unmasked areas. Thus, the total loss is defined as: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{ldm}}+\mathcal{L}_{\text{attn}}+\lambda_{\text{info}}(\mathcal{L}_{\text{CLS}}+\lambda_{\text{seg}}\mathcal{L}_{\text{SEG}})+\lambda_{\text{bg}}\mathcal{L}_{\text{BG}},$ with the following loss weights: $\lambda_{\text{info}}=0.05$, $\lambda_{\text{seg}}=10.0$, $\lambda_{\text{bg}}=0.01$.